2). Explain the difference between change in supply and change in quantity supplied Use graph to support your answer I 3a
Answer #1

As per policy we have to answer first question, I have answered more than one

1) A change in supply causes a shift in the entire supply curve in response to something besides price such as production costs, taxes, changes in input prices, changes in regulations, and improvement in technology as enclosed in graph-1. A change in quantity supplied causes a movement along the supply curve due to the change in price as enclosed in graph-2

2) Percentage change in Demand = [(Demand End - Demand Start) / Demand Start] = (94 - 90) / 90 = 0.04444

Percentage change in Income = [(Income End - Income Start) / Income Start] = (45,000 - 44,000) / 44,000 = 0.027273

Income Elasticity of Demand = Percentage change in Demand / Percentage change in Income

= 0.04444 / 0.027273 = 1.96

As income elasticity of demand is above 1 thus the good is a normal product and a luxury

GRAPH-1 GRAPH-2 Price S1 SO S2 Price Price So Quantity Quantity